OpenStudy (anonymous):

simplify sqrt 18x^9y^4

jagr2713 (jagr2713):

3x^4y^2√2x

OpenStudy (anonymous):

\[\sqrt{18x ^{9}y ^{4}}\]

OpenStudy (anonymous):

Thanks for being so fast.

jagr2713 (jagr2713):

yw

OpenStudy (anonymous):

could you help me with\[\sqrt{2}-\sqrt{10} / \sqrt{2}+\sqrt{10}\]

jagr2713 (jagr2713):

1/2(−3+√5)